# Malsight: Exploring Malicious Source Code and Benign Pseudocode for Iterative Binary Malware Summarization

### v1

@@ 3.3.1 Dataset For Code Summary Model
In this paper, our key insight is that the code summary model requires two capabilities, understanding of Malware Functionality and adaptability to disassembled pseudocode formats (including the ability to deal with annotated code). Therefore, We build source-based malware datasets and pseudo-code-based benign software datasets to train the model on these two capabilities separately. Based on Sourcefinder [49], we were able to find malware source repositories from GitHub. We generate descriptive labels for the extracted malware functions using a sophisticated language model, followed by manual verification and optimization. Meanwhile, we use the Capybara dataset provided by BinT5 [19] (a benign software dataset) to train the model adaptability to pseudocode structure.
3.3.2 Dataset For Evaluation Model
@@
In the evaluation phase, the evaluation method is used to measure the similarity between the model generation results and the reference results to evaluate the quality of the model generation. Given two sentences (generation results and the reference results) $S_{g}$ and $S_{r}$, most evaluation methods output a score $Score$ as the evaluation result. Therefore, if considering the use of machine learning methods, it is necessary to construct a dataset in $\{S_{g},S_{r},Score\}({Score}\in[0,1])$ format. The challenge is that when a dataset of $\{S_{g},S_{r}\}$ is obtained, it is a difficult job to obtain an accurate $Score$. In our subsequent work, we propose a reasonable algorithmic flow for constructing labeled datasets $EvaS$.
3.3.3 Dataset For Static Annotater
@@
In annotation generation process, the static annotater includes a core information extraction module (described in detail in Section 4.2.1). Due to the difficulty in accurately completing the required functions using static methods, we use a machine learning model to complete the sequence labeling task of the pseudocode. By constructing the dataset AnnoS, we have constructed the dataset required for model training and testing.
TABLE I: The Proposed Datasets
@@
AnnoS 95,000 pseudo 3.87 Train & Test
@@
To sum up, we mainly completed the construction of three sets of datasets in different application fields, as shown in Table I.
4 Code Summarization Workflow
Following our breakdown of the malware code summary task in Figure 2, our implementation first extracts the reverse function list, and then sequentially generates static and dynamic annotations for the items in the function list, and finally passes them into the code summary model.
In this process, we completed the training and designing of two Domain-Specific large model, the applying of a General-Purpose large model and the implementation of several algorithms. we cover the implementation of each step separately in this section.
4.1 Function List Extraction
As mentioned earlier in Section 3.1, in the first step of the workflow, we extract the list of reverse functions from the CFG of the malware binary.
Since the existing methods [50] do not give a completely accurate CFG extraction flow, we implement a pluggable CFG extraction module. It is used to provide us with a processing scheme from binary file $B_{Mal}$ to digraph $G_{M}$ as the CFG. By an inverse topological traversal algorithm, it is extracted from CFG in the opposite direction of the call chain, expressed as $L_{G_{M}}=[f_{1},f_{2},...,f_{n}](n=G_{M}.vertices)$, Where $f_{i}$ represents the $i$ th function of inverse topological order of $G_{M}$.
In this study, algorithm $REsort$ was constructed, expressed as $L_{G_{M}}=REsort(G_{M})$. (See Appendix B for details) By applying algorithms such as Tarjan [51], Dijkstra [52], and Depth First Search (DFS), this approach successfully addressed the obstacles caused by cyclic calls and partially connected graphs in the process of generating reverse order lists. Function columns $L_{G_{M}}$follow the order from outside to inside in the function call diagram to ensure that the later-called function is first in the function list and is processed first in subsequent steps.
The following process takes the function instances from $L_{G_{M}}$ in a forward order to achieve the order of restoration from the outer layer of the CFG diagram to the inner layer, specifically, recovering from the outer API call to the main function.
4.2 Annotation Generation
@@
In the order of traversal provided by the reverse function list, annotations are added to each function in turn to provide richer information. Annotations can be divided into static and dynamic types.
We have observed that attackers frequently employ techniques like stripping to hinder reverse engineering efforts. This process removes crucial symbolic information, such as identifier names, from the binaries, leading to significant semantic loss in the generated pseudocode. As a result, code summary models for malware face substantial challenges, and the performance of code summarization tasks is adversely affected.
To address this issue, we propose utilizing dynamic and static annotation to supplement the semantic information and enhance the features of the stripped pseudocode as Figure 4 shows. This approach aims to compensate for the poor transferability caused by the stark dissimilarity between the source code and the stripped decompiled code.
4.2.1 Static Annotation
Based on our long-term exploration of pseudocode for stripped malware, we deem that although the stripped pseudocode has a serious semantic loss, some extremely critical API calls (such as operating system APIs) and some special forms of strings that are preserved after stripping provide us with ideas for behavior analysis and semantic recovery of malware.
Consequently, we consider building a static annotation module to provide additional information for subsequent code summaries. In general, the static annotation process can be divided into three parts: sequence labeling, online retrieval and annotation generation.
Sequence labeling model: By manually labeling approximately 300,000 tokens within nearly 80,000 functions, we construct a labeled dataset $Set_{SL}$ based on $BenignS$ to train the sequence labeling model.
@@ 4.2.1 Static Annotation
$Set_{CSL}=\{d_{0},d_{1},...,d_{N}\}$ (1)
Subsequently, we opted to utilize the CodeBERT [39] model for training the sequence labeling task using this dataset CSL.
$B$ is the CodeBERT base model, and $C$ is the classifier head. As a sequence is inputted, the complete model outputs the predicted labels of its tokens. Equation (2) and (3) formalize this.
$B(s_{i})=\{o_{0},o_{1},...,o_{n-1}\}\vspace{-0.2cm}$ (2)
@@
We further formalize the target function in equation (4), where $l_{i_{c}}$ is the truth label and $y_{i_{c}}$ is the Softmax probability for the $c^{th}$ class.
@@
Label to annotation: Once we can get the key API calls and special strings in the stripped pseudocode, we use GitHub Code Search [53] to retrieve the relevant context in the GitHub repositories. In the implementation, we keep the first three blocks of the search results (this is because the code search has already sorted the relevance of the results [54]).
The outcome of random sampling and manual discrimination reveals that approximately 54.8% of the function contexts contain code comments closely associated with the functionality of the function. For the remaining functions, nearly 90% also offer contextual information related to the function’s operation, such as parameter names, interconnected functions, and processing logic. Only a small number of functions yield invalid search results.
The filtered and preprocessed code snippets will be continuously input into the prompt-based generic model for generating static annotation.
4.2.2 Dynamic Annotation
As shown in Figure 4 (the blue parts represent the steps in which the annotation was added), the summary of the callee is provided to the caller as a complement to the semantic information, which we define as dynamic annotation. This is consistent with the actual analysis flow of binary malware analysis by reverse workers, i.e., analyzing the call relationship from the inner layer of the CFG diagram to the outer layer (corresponding function list generated in Section 4.1) and summarizing the function from the outside in (corresponding to the passing of dynamic annotation). In this way, we can make full use of the dynamic behavior characteristics implied by the call relationships between functions in the pseudocode.
4.3 Building Malware Datasets
@@
Out of 2,289 GitHub repositories that were determined to be malware, we extracted close to 30K functions. We filter for functions that repeat, shorter than five lines, and format-challenged functions, resulting in a dataset of 89,609 functions. (The lack of strict filtering may lead to overlap between the train and test sets, consequently yielding inflated results.)
Unlike benign open-source projects that are well maintained, most of the functions we extract do not have comments in context for us to label as code summaries. Fortunately, the semantic information in the source code is rich enough that we used a well-designed prompt to complete the code summary for us via GPT3.5-Turbo [56]. In this way, we extract dataset MalS in equation (6) format. Further, we build a 500-function dataset MalP for testing our model. The reason why dataset MalP has a relatively small quantity is that it was obtained by manually compiling and decompiling the git repository. MalP is compiled from the makefile provided by the developer, so we have not mentioned the configuration related to compilation optimization.
@@
As mentioned above, the malware training set we built was made up of source code, so another dataset was needed to help our model understand the stripped function features. In the transfer learning option, we used the Capybara dataset (a benign software dataset) provided by BinT5 [19] and annotated it to provide our model with adaptations to the annotated code summary task. It is provided as $Set_{Capybara}$, and we process it as $Set_{benignC}$, where the $ANN()$ is the static annotation generation method to form PL-NL bimodal data $f_{n}^{ann}$.
@@
In the process of training the code summary model, we use $Set_{MalS}$ and $Set_{benignC}$ to complete the transfer learning process, and $Set_{MalP}$ to test during the evaluation phase.
@@
Figure 5: Code Summary Topic Distribution. The results of topic analysis for ${Set_{MalS}}$ indicate that its distribution conforms to the functional distribution of malware functions.
@@
For $Set_{MalS}$, 25% of the samples were extracted for thematic analysis of function functionality, as shown in Figure 5. Our analysis results confirm that $Set_{MalS}$ and $Set_{BenignC}$ have significantly different theme distributions. Among a large number of security-related functions in the former, there are an unignorable number of codes for lock, permissions, and encryption, while in the latter, there are a large number of codes related to game logic and driver calls which irrelevant to malware.
4.4 Code Summary Model
@@
In Malsight, we fine-tune the CodeT5+ model to enable transfer learning from source code summarization to decompiled pseudocode summarization. To achieve this, we have designed two distinct phases of fine-tuning to facilitate smooth transfers that accommodate variations in data characteristics and distribution biases across different datasets.
Phase 1. In the initial phase, we fine-tune the model that uses the dataset we built consisting of malware source code $Set_{MalS}$. The source code of malware has a similar semantic structure to the code in ordinary scenarios, but it also has behavioral or semantic features that the latter does not exist, which are represented by some code fragments with malicious purposes, such as self-replication and propagation, illegal access to system resources, and vulnerability exploitation. Through this phase of fine-tuning, the model can learn the behavioral or semantic features of these malicious codes. In this phase, the poor effects of data shift are addressed by using $Set_{MalS}$ with the same functional distribution as malware in the real world.
Phase 2. In this phase, we ask the model to learn the corresponding semantic information from the pseudo-code and the annotation text simultaneously to better assist the model in generating high-quality code summaries. We use the PL-NL bimodal dataset BenignC to fine-tune the encoder of the model which we fine-tuned in the previous phase in case the decoder’s parameters are frozen. By only fine-tuning the encoder, we allow the model to adjust itself to better suit the specific task at hand, without updating the weights too much. This can not only lead to a more robust and generalized understanding of the data but also shorten the training time due to the reduction in parameters. Especially, the raw input is processed into a standardized format as equation (10) where $t_{c_{i}}$ is the tokens from a code sequence, $t_{a_{i}}$ is the tokens from the corresponding annotation, $t_{sep}$ is a special token in CodeT5+ to separate the inputs in different modes. Inputs in such a format can assist the encoder in distinguishing the difference between two modes and learning the correlation between them.
$f_{n}^{ann}\to\{t_{c_{0}},...,t_{c_{n-1}},t_{sep},t_{a_{0}},...,t_{a_{m-1}}\}$ (10)
After the above two phases of fine-tuning, the model basically has the ability to accept bimodal input composed of pseudocode and annotation text and generate code summary. We use this model to summarize the entire malware in reverse of the function call order.
5 Evaluation Method
@@
As mentioned in Section 3.3.2, the evaluation algorithm for code summary tasks should accept the reference as input and separate available and unavailable generated results. In this section, we introduce our exploration of code summary dataset construction (EvaS) and evaluation model construction (BLEURT-sum) respectively.
5.1 Evaluation Dataset Construction

### v3

@@ III-C1 Dataset For Code Summary Model
Our critical insight is that the code summary model requires two capabilities, understanding of Malware Functionality and adaptability to disassembled pseudocode formats (including the ability to deal with annotated code). Therefore, we separately build datasets from malware source code obtained from GitHub [50] and benign pseudocode from the Capybara [20] dataset. They correspond to MalS and BenignC, respectively, where labels are generated through a carefully designed annotation process that combines large language models with human verification.
We built the MalP dataset by compiling and decompiling 20 randomly selected repositories from the above Github malicious code repositories, resulting in 500 functions written in pseudocode for evaluation purpose.
III-C2 Dataset For Evaluation Model
@@
In the evaluation phase, the evaluation method is used to measure the similarity between the model generation results and the reference results to evaluate the quality of the model generation. Given two sentences (generation results and the reference results) $S_{g}$ and $S_{r}$, most evaluation methods output $Score$ as the evaluation result. Given the tremendous potential shown by machine learning methods, it is necessary to construct a dataset in the format ${S_{g},S_{r},Score}({Score}\in[0,1])$ when considering their use. The challenge lies in obtaining an accurate $Score$ once a dataset of ${S_{g},S_{r}}$ is constructed. In our subsequent work, we propose a reasonable algorithmic flow for constructing labeled datasets $EvaS$.
III-C3 Dataset For Static Annotater
@@
In the annotation generation process, the static annotater includes a core information extraction module (detail in Section IV-B1). Due to the difficulty in accurately completing the required functions using static methods, we use a machine learning model to complete the sequence labeling(SL) task of the pseudocode. By constructing the dataset CSL, we have constructed the dataset required for model training and testing.
TABLE I: The Three Proposed Datasets
@@
CSL 95,000 pseudo 3.87 Train & Test
@@
To sum up, we construct three datasets to address the different tasks arising from the various stages of the Malsight framework, as shown in Table I.
IV Code Summarization Workflow
Following our breakdown of the malware code summary task in Fig. 2, we construct comprehensive systems incorporating multiple LLMs and algorithms.
IV-A Function List Extraction
In the first module of Malsight, we extract a reverse function list from the malware’s FCG. Since existing methods [51] don’t fully ensure accurate FCG extraction, we implement a pluggable module to map the binary $B_{Mal}$ to the FCG digraph $G_{M}$. Using an inverse topological traversal algorithm, $REsort$, we generate the reverse order list $L_{G_{M}}$, where functions are arranged according to their inverse topological order.
Algorithm 1 REsort
1:The FCG graph $G_{M}$ for a malware binary file.
2:The reverse topsort list $L_{G}$ of the function call graph.
3:function REsort($G_{M}$)
4:     for each vertex $v$ in $G_{M}.vertices$ do
5:         if $!v.seen$ then
6:              new $G_{M}^{\prime}$
7:              $G_{M}^{\prime}.vertices\leftarrow$ call Tarjan($v$, $G$)
8:         end if
9:     end for
10:     $G_{M}^{\prime}\leftarrow$ call BuildTarGraph($G_{M}^{\prime}.vertices$, $G_{M}$)
11:     $L_{G_{M}^{\prime}}\leftarrow$ call RetopSort($G_{M}^{\prime}$)
12:     $dist[]\leftarrow$ call Dijkstra($G_{M}$)
13:     $L_{G_{M}}\leftarrow[]$
14:     for each vertex $v^{\prime}$ in $L_{G_{M}^{\prime}}$ do
15:         $mindist\leftarrow\infty$, $idx\leftarrow-1$
16:         for each $ver$ in $v^{\prime}.subvertex$ do
17:              if $dist[ver]<mindist$ then
18:                  $mindist\leftarrow dist[ver]$, $idx\leftarrow ver$
19:              end if
20:         end for
21:         $L_{G_{M}}$.append(DFS($idx$, $v^{\prime}$))
22:     end for
23:end function
Algorithm 1 shows that equipped with Tarjan, Dijkstra, and DFS, our proposed algorithm is able to address challenging issues such as cyclic calls and partially connected graphs. The function list $L_{G_{M}}$ is processed from outer API calls to the main function to ensure the correct restoration order. Specifically, $L_{G_{M}}$ follows the call hierarchy from the outermost to the innermost layers, ensuring that functions called later appear earlier in the list and are processed first in subsequent steps.
IV-B Annotation Generation
@@
In $L_{G_{M}}=\{f_{1},f_{2}...f_{len(L_{G_{M}})}\}$, we sequentially add static and dynamic annotations to enrich the information for each function $f_{i}$.
As mentioned, attackers often use techniques like stripping to remove symbolic information, leading to significant semantic loss in the pseudocode and hindering code summarization models. To address this, we propose using dynamic and static annotations to supplement the missing semantics and enhance pseudocode features. This approach, as illustrated in Fig.  4, helps bridge the gap between source code and stripped pseudocode.
IV-B1 Static Annotation
Through our extensive study of stripped malware pseudocode, we identified that critical API calls (e.g., OS APIs) and certain preserved strings offer valuable clues for malware behavior analysis and semantic recovery despite the significant semantic loss. Thus, we design a static annotation module to enhance code summaries, which operates in three key steps: sequence labeling, online retrieval, and annotation generation.
Sequence labeling model: By manually labeling approximately 300,000 tokens within nearly 95,000 functions with the assistance of lex [52], we construct a labeled dataset CSL based on BenignC to train the sequence labeling model.
@@ IV-B1 Static Annotation
$CSL=\{d_{0},d_{1},...,d_{N}\}$ (1)
Subsequently, we utilize the CodeBERT [39] model for training the SL task using this CSL. As shown in equation (2) and (3), $B$ is the CodeBERT base model, and $C$ is the classifier head. As a sequence $s_{i}$ is inputted, the complete model outputs the predicted labels $y_{i}$ of its tokens $o_{i}$.
$B(s_{i})=\{o_{0},o_{1},...,o_{n-1}\}\vspace{-0.1cm}$ (2)
@@
Further, the target function is formalized in equation (4), where $l_{i_{c}}$ is the truth label and $y_{i_{c}}$ is the Softmax probability for the $c^{th}$ class.
@@
Label to annotation:
Once we obtain the key API calls and special string literals from the stripped pseudocode, we utilize GitHub Code Search [53] to retrieve relevant contextual information from public GitHub repositories.
In practice, we construct a query based on the extracted identifiers and submit it to the code search engine. To ensure relevance while maintaining efficiency, we retain only the top three retrieved blocks for each query. This decision is motivated by the observation that GitHub Code Search sorts results by estimated relevance according to their ranking algorithm [54], and that the top-ranked blocks typically contain the most pertinent information.
We conducted random sampling and manual inspection across the retrieved results. The analysis shows that approximately 54.8% of the retrieved function contexts contain comments explicitly describing the function’s intended functionality. Among the remaining functions, nearly 90% still provide valuable contextual signals, such as meaningful parameter names, semantically related helper functions, control flow hints, and call dependencies. Only a small minority of cases yield irrelevant or unusable retrievals, often due to noise in the search index or ambiguous identifiers. After retrieval, the code snippets undergo a filtering and preprocessing step. Specifically, we discard results that are excessively short, semantically incomplete, or duplicate the input pseudocode without adding new context. The remaining high-quality snippets are then incrementally injected into the prompt of a generic, instruction-following large language model to generate static natural language annotations for each function, denoted by $AnnoS(f_{i})$. This retrieval-augmented annotation strategy significantly enriches the available context for static analysis and serves as an important intermediate process in our overall framework.
IV-B2 Dynamic Annotation
As shown in Fig. 4 (the blue parts represent the steps in which the annotation was added), the summary of the callee is provided to the caller as a complement to the semantic information, which we define as dynamic annotation. This is consistent with the actual analysis flow of binary malware analysis by reverse workers, i.e., analyzing the call relationship from the inner layer of the FCG diagram to the outer layer (corresponding function list generated in Section IV-A) and summarizing the function from the outside in (corresponding to the passing of dynamic annotation). In this way, we can make full use of the dynamic behavior characteristics implied by the call relationships between functions in the pseudocode, denoted by $AnnoD(f_{i})$. The function $f_{i}=f_{i}\oplus AnnoS(f_{i})\oplus AnnoD(f_{i})$ will be passed into the code summary model for later processing.
IV-C Building Malware Datasets
@@
Out of 2,289 GitHub repositories that are determined to be malware, we extract close to 30K functions. We filter out functions that are repetitive, shorter than five lines, or have formatting issues, resulting in a dataset of 89,609 functions. (The lack of strict filtering may lead to overlap between the train and test sets, consequently yielding inflated results.)
Unlike well-maintained open-source projects, most extracted functions lack comments for use as code summaries. However, the source code’s rich semantics allowed us to generate summaries using a well-designed prompt with LLM, and engaged five human experts to assist in reviewing and refining the generated summaries, resulting in the MalS dataset in equation (6). We built the MalP dataset with 500 functions, selected from 20 randomly chosen repositories. Its smaller size is due to the difficulty of manually compiling and decompiling poorly maintained malware repositories, from which we picked functions with strong functionality characteristics.
@@
Since MalS is a source code training set, we need an additional dataset to help our model understand stripped pseudocode features. The Capybara dataset provided by BinT5 [20] is a benign pseudocode dataset. We process the Capybara dataset ($Set_{Capybara}$) using our system’s static annotation module ($ANNS()$), transforming it into BenignC.
@@
When training the code summary model, we use MalS and BenignC to complete the transfer learning tuning process, and $Set_{MalP}$ to test during the evaluation phase.
@@
Figure 5: Topic distribution of code summary. The results of topic analysis for ${Set_{MalS}}$ indicate that its distribution conforms to the functional distribution of malware functions.
@@
Additionally, 25% samples from MalS were randomly extracted for thematic analysis of function functionality, as shown in Fig. 5. Our analysis results confirm that MalS and BenignC have significantly different theme distributions. Among a large number of security-related functions in the former, there are an unignorable number of codes for lock, permissions, and encryption, while in the latter, there are a large number of codes related to game logic and driver calls which irrelevant to malware.
IV-D Code Summary Model
@@
In Malsight, we fine-tune the CodeT5+ model to enable transfer learning from source code summarization to decompiled pseudocode summarization. To achieve this, we design two distinct phases of fine-tuning to facilitate smooth transfers that accommodate variations in data characteristics and distribution biases across different datasets.
In Malsight, we leverage transfer learning to shift the model from source code summarization to decompiled pseudocode summarization. This is achieved by utilizing two datasets: MalS (source code dataset) and BenignC (pseudocode dataset), aiming to approximate the ideal training effect on an unachievable dataset $Set_{ideal}$.
First, the model is trained on MalS for semantic features of the malware source code. For each sample $(f_{n}^{Sou},f_{n}^{sum})$ in MalS, where $f_{n}^{Sou}$ represents a source code snippet and $f_{n}^{sum}$ represents the corresponding summary, the model minimizes the loss function as equation (10). Here, $M_{\theta}$ denotes the model parameters, and $\mathcal{L}$ represents the loss function (e.g., cross-entropy loss). The output of this stage is the encoded feature representation of the source code $h_{MalS}=\text{Encoder}_{\theta}(x_{i})$.
$L_{MalS}=\frac{1}{n_{MalS}}\sum_{i=1}^{n_{MalS}}\mathcal{L}(M_{\theta}(f_{n}^{Sou}),f_{n}^{sum})$ (10)
Next, the model undergoes transfer learning using BenignC (pseudocode dataset) to further refine its capabilities. For each sample $(f_{i}^{ann},f_{n}^{sum})$ in BenignC, where $f_{i}^{ann}$ is the pseudocode sequence and $f_{n}^{sum}$ is the annotation, the data processing pipeline is shown in equation (11).
$f_{i}^{ann}\to\{fc_{0}^{ann},\dots,fc_{p-1}^{ann},f_{sep},fc_{0}^{ann},\dots,fa_{q-1}\}$ (11)
The input pseudocode and annotation are concatenated into a bimodal sequence where $fc_{p}$ represents tokens from the pseudocode, $fa_{q}$ represents tokens from the annotation, and $f_{sep}$ is a special separator token.
The T5 model’s encoder processes this sequence of pseudocode and annotations, generating an encoded representation $h_{benignC}=\text{Encoder}_{\theta}(f_{i}^{ann})$. This representation serves as the basis for generating the corresponding code summary. To optimize the model’s performance, we minimize the loss function defined in equation (12), where $M_{\theta}$ represents the parameters of the model, $f_{i}^{ann}$ is the input pseudocode sequence, and $f_{i}^{sum}$ is the target summary.
$L_{benignC}=\frac{1}{n_{benignC}}\sum_{i=1}^{n_{benignC}}\mathcal{L}(M_{\theta}(f_{i}^{ann}),f_{i}^{sum})$ (12)
This loss function captures the difference between the predicted and target summaries across the entire dataset. The model leverages transfer learning by training on both MalS (source code) and BenignC (pseudocode), allowing it to learn generalized features from source code and apply them to pseudocode summarization. By facilitating knowledge transfer between these datasets, the T5 encoder ensures consistency and improves its ability to handle different data modalities, enhancing performance on both source code and pseudocode tasks.
$L_{total}=\alpha L_{MalS}+\beta L_{benignC}$ (13)
The overall objective loss for the transfer learning process is defined in equation (13). Specifically, $\alpha$ and $\beta$ are balancing factors between the two learning objectives. This method allows the model to adapt to the semantic structure of pseudocode and generate high-quality code summaries.
V Evaluation Method
@@
As mentioned in Section III-C2, the evaluation algorithm for code summary tasks takes the reference as input and distinguish between valid and invalid generated results. This section introduces our exploration of EvaS and BLEURT-sum, respectively.
V-A Evaluation Dataset Construction